CLEAR Weekly Operation Meeting

Europe/Zurich
CERN

CERN

Videoconference
CLEAR Weekly Operation Meeting
Zoom Meeting ID
67316811109
Host
Roberto Corsini
Useful links
Join via phone
Zoom URL

MINUTES

Actions from previous week:

  • Check the availability of a beam pipe to allow the installation of only two CLIC cavity BPM      > ongoing
  • Further look to solve the noise in the plasma lens pump.    > ongoing
  • Devise a way of doing laser alignment on the in-air test stand. > done, to be completed for spectrometer line
  • Understand where's the end of the trigger cables in the klystron gallery. > to be completed
  • Look at moving one of the unused CTF3 front ends to the barrack to aid data collection. > to be verified
  • Reconnection of the thermalized water pipes on the X-band ACS     > ongoing
  • Re-installation of safety valve on gas pipe for P.L.    > ongoing, long term
There are minutes attached to this event. Show them.
    • 9:00 AM 9:15 AM
      Report of last week operation 15m
    • 9:15 AM 9:35 AM
      Week operation program 20m
    • 9:35 AM 9:55 AM
      AOB 20m